what is the y coordinate of the vertex of a parabola wit the following equation?
y=x^2-6x+14

Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(7384)

y=x^2-6x+14
.
The 'x' coordinate of the vertex can be found by:
x = -b/(2a)
x = -(-6)/(2(1))
x = 6/2
x = 3
.
To find the 'y' coordinate, plug it back into the original equation:
y=3^2-6(3)+14
y=9-18+14
y = 5 (this is your answer)

Answer by solver91311(24713)



I'll do you one better. I'll tell you how to find the -coordinate of the vertex of any parabola with equation in standard form.

Given



first find the -coordinate of the vertex using:



then evaluate your function for that value of the independent variable:




John

My calculator said it, I believe it, that settles it